# YITP-25-129, KOBE-COSMO-25-16

### v1

@@ 4 Nonclassicality of gravitational waves
$\beta^{(+)}_{kk^{\prime}}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{3i}{64\pi^{3}M_{\rm p}^{2}}\mu(a\Omega)^{4}$ (4.4)
@@
$\hskip-42.67912pt\times\frac{e^{i(\omega_{k}+\omega_{k^{\prime}})t^{\prime}}}{\sqrt{\omega_{k}\omega_{k^{\prime}}}}\left[\frac{\sin^{2}\theta}{2}+\frac{1+\cos^{2}\theta}{2}\cos(2\Omega t^{\prime}-2\varphi)\right]\left[\frac{\sin^{2}\theta^{\prime}}{2}+\frac{1+\cos^{2}\theta^{\prime}}{2}\cos(2\Omega t^{\prime}-2\varphi^{\prime})\right]\$
@@
$\beta^{(\times)}_{kk^{\prime}}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{3i}{64\pi^{3}M_{\rm p}^{2}}\mu(a\Omega)^{4}$ (4.5)
@@
Constraints on the squeeing parameters have been discussed in the literature [21], based on LIGO data [22, 22].
We now evaluate the squeezing parameter at a location far from the binary system. The expressions derived above correspond to the values at the source. In this paper, we assume that no decoherence processes act on the gravitational waves during their propagation to the observer. Under this assumption, the value of the squeezing parameter at the detector is identical to that at the source. From Eqs. (4.4) and (4.5), we see that the squeezing parameter attains its largest value is obtained in the case where ${\bm{k}}=-{\bm{k}^{\prime}}$ and $\omega_{k}=\omega_{k^{\prime}}=k=2\Omega$. The squeezing parameter at the location of the binary system can be estimated as
In the second expression, we have used the relation $\omega=2\pi f$. As in Section 3, we adopt the parameters of GW150914 [1]. In this case, we obtain
$\zeta\simeq 2\times 10^{-3}\left(\frac{\mu}{16~M_{\odot}}\right)\left(\frac{a\Omega}{0.41}\right)^{4}\left(\frac{f}{68~{\rm Hz}}\right)\ .$ (4.7)
Thus, the quantum state of gravitational waves from GW150914 is characterized by a squeezing parameter of order $10^{-3}$. While this estimate is based on the ISCO, the squeezing is expected to be even stronger just before the merger. For example, in GW150914 the frequency of the gravitational wave reached up to $150~{\rm Hz}$.
5 Conclusion
In this work, we investigated how astrophysical binary black holes generate a quantum state of gravitons. We modeled the binary system as a classical source. As is well known, when a classical source couples linearly to a quantum field, it produces a coherent state.
We confirmed that classical gravitational waves can be described within the coherent state framework. We then extended the analysis to the next-order effects and showed that binary black holes can also generate squeezed states. Furthermore, we estimated the degree of squeezing and found that, for GW150914, the squeezing parameter is of order $10^{-3}$. This result indicates a measurable deviation from a purely coherent state. Consequently, there may be a realistic opportunity to probe the nonclassicality of gravitational waves with current and future detectors such as LIGO, Virgo, and KAGRA.
Since we have obtained a squeezed coherent state, Hanbury–Brown–Twiss interferometry may be employed to probe the nonclassicality of the graviton state [23, 24, 25] with new technology [26, 27]. It would also be interesting to extend our analysis to other sources of gravitational waves. In this paper, we have demonstrated how a quantum state is generated by binary black holes under the assumption of the standard vacuum. However, if primordial gravitational waves generated during inflation are taken into account, an additional enhancement of squeezing is expected. Thus, by observing the squeezing of gravitational waves from binary black holes, one may obtain indirect information about the early universe. The detail will be reported separately [28].
Acknowledgments

### v2

@@ Appendix A Derivation of the coherent parameter
$\alpha^{(P)}(\bm{k})\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{i}{(2\pi)^{3/2}}\sum_{N=1,2}\int^{t}dt^{\prime}\frac{\gamma_{N}^{3}m_{N}}{M_{\rm p}}\frac{e^{i\omega_{\bm{k}}t^{\prime}}}{\sqrt{2\omega_{\bm{k}}}}e^{(P)}_{ij}(\bm{k})v_{N}^{i}v_{N}^{j}e^{-i\bm{k}\cdot\bar{\bm{x}}_{N}}\ .$ (A.1)
Choosing the $z$-axis in $\bm{k}$} space to align with the position vector $\bm{x}$}, we parametrize the wave vector as $\bm{k}=k(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta)$. The coherent state parameter is specifically written as
$\alpha^{(+)}(\bm{k})\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{i}{(2\pi)^{3/2}}\frac{\mu(a\Omega)^{2}}{\sqrt{2}M_{\rm p}}\int^{t}dt^{\prime}\frac{e^{i\omega_{\bm{k}}t^{\prime}}}{\sqrt{2\omega_{\bm{k}}}}\left(\frac{\sin^{2}\theta}{2}+\frac{1+\cos^{2}\theta}{2}\cos(2\Omega t^{\prime}-2\varphi)\right)$ (A.2)
$\alpha^{(\times)}(\bm{k})\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{i}{(2\pi)^{3/2}}\frac{\mu(a\Omega)^{2}}{\sqrt{2}M_{\rm p}}\int^{t}dt^{\prime}\frac{e^{i\omega_{\bm{k}}t^{\prime}}}{\sqrt{2\omega_{\bm{k}}}}\cos\theta\sin(2\Omega t^{\prime}-2\varphi)$ (A.3)
where
$k_{x}x_{1}+k_{y}y_{1}\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{m_{2}}{M}ak(\sin\theta\cos\varphi\cos(\Omega t^{\prime})+\sin\theta\sin\varphi\sin(\Omega t^{\prime}))\ ,$ (A.4)
$k_{x}x_{2}+k_{y}y_{2}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{m_{1}}{M}ak(\sin\theta\cos\varphi\cos(\Omega t^{\prime})+\sin\theta\sin\varphi\sin(\Omega t^{\prime}))\ .$ (A.5)
We also used the explicit form of polarization tensors
$e^{(+)}_{ij}\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{1}{\sqrt{2}}\begin{pmatrix}\cos^{2}\theta\cos^{2}\varphi-\sin^{2}\varphi&(1+\cos^{2}\theta)\sin\varphi\cos\varphi&-\displaystyle\frac{1}{2}\sin 2\theta\cos\varphi\\[12.0pt] (1+\cos^{2}\theta)\sin\varphi\cos\varphi&\cos^{2}\theta\sin^{2}\varphi-\cos^{2}\varphi&-\displaystyle\frac{1}{2}\sin 2\theta\sin\varphi\\[12.0pt] -\displaystyle\frac{1}{2}\sin 2\theta\cos\varphi&-\displaystyle\frac{1}{2}\sin 2\theta\sin\varphi&\sin^{2}\theta\end{pmatrix}$, (A.6)
$e^{(\times)}_{ij}\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{1}{\sqrt{2}}\begin{pmatrix}-\cos\theta\sin 2\varphi&\cos\theta\cos 2\varphi&\sin\theta\sin\varphi\\[12.0pt] \cos\theta\cos 2\varphi&\cos\theta\sin 2\varphi&-\sin\theta\cos\varphi\\[12.0pt] \sin\theta\sin\varphi&-\sin\theta\cos\varphi&0\end{pmatrix}\$. (A.7)
The expectation value of the metric operator in the coherent state is
[…]
As an example, we show the calculation of the $h_{xx}$ component.
[…]
where we defined $F^{(+)}(t,r)$ and $F^{(\times)}(t,r)$ as follows.
$F^{(+)}(t,r)\hskip-5.0pt$ $\equiv$ $\hskip-5.0pt\int_{0}^{\infty}rdk~\frac{1}{2}\int_{0}^{\pi}\sin\theta d\theta~\frac{1}{2\pi}\int_{0}^{2\pi}d\varphi~\frac{1}{2\pi}\int_{-\infty}^{t}kdt^{\prime}$ (A.10)
$\times\left\{i\left[\gamma^{3}_{1}\left(\frac{m_{2}}{M}\right)e^{-i(k_{x}x_{1}+k_{y}y_{1})}+\gamma^{3}_{2}\left(\frac{m_{1}}{M}\right)e^{-i(k_{x}x_{2}+k_{y}y_{2})}\right]e^{ikr\cos\theta-i\omega_{k}t}e^{i\omega_{k}t^{\prime}}\right$.
$\hskip 24.0pt\left.\times e_{xx}^{(+)}\left(\frac{\sin^{2}\theta}{2}+\frac{1+\cos^{2}\theta}{2}\cos(2\Omega t^{\prime}-2\varphi)\right)+{\rm c.c.}\right\}\,$,
$F^{(\times)}(t,r)\hskip-5.0pt$ $\equiv$ $\hskip-5.0pt\int_{0}^{\infty}rdk~\frac{1}{2}\int_{0}^{\pi}\sin\theta d\theta~\frac{1}{2\pi}\int_{0}^{2\pi}d\varphi~\frac{1}{2\pi}\int_{-\infty}^{t}kdt^{\prime}$ (A.11)
$\times\left\{i\left[\gamma^{3}_{1}\left(\frac{m_{2}}{M}\right)e^{-i(k_{x}x_{1}+k_{y}y_{1})}+\gamma^{3}_{2}\left(\frac{m_{1}}{M}\right)e^{-i(k_{x}x_{2}+k_{y}y_{2})}\right]e^{ikr\cos\theta-i\omega_{k}t}e^{i\omega_{k}t^{\prime}}\right$.
$\hskip 24.0pt\left.\times e_{xx}^{(\times)}\cos\theta\sin\left(2\Omega t^{\prime}-2\varphi\right)+{\rm c.c.}\right\}\ ,$
where $r=|\bm{x}|$ and the integrals can be evaluated approximately. Next we assume that a binary black hole system with two masses of approximately equal, and set
$\gamma^{3}_{1}\left(\frac{m_{2}}{M}\right)e^{-i(k_{x}x_{1}+k_{y}y_{1})}+\gamma^{3}_{2}\left(\frac{m_{1}}{M}\right)e^{-i(k_{x}x_{2}+k_{y}y_{2})}\simeq 1\ .$ (A.12)
Under this assumption, we obtain
$\hskip-12.0ptF^{(+)}(t,r)$
$=$ $\hskip-5.0pt\int_{0}^{\infty}rdk~\frac{1}{2}\int_{0}^{\pi}\sin\theta d\theta~\frac{1}{2\pi}\int_{0}^{2\pi}d\varphi~\frac{1}{2\pi}\int_{-\infty}^{t}kdt^{\prime}$
$\left\{\frac{i}{2}e^{ikr\cos\theta-i\omega_{k}t}\left[e_{xx}^{(+)}\sin^{2}\theta e^{i\omega_{k}t^{\prime}}+e_{xx}^{(+)}\frac{1+\cos^{2}\theta}{2}e^{i\omega_{k}t^{\prime}}\left(e^{i2\Omega t^{\prime}}e^{-2i\varphi}+e^{-i2\Omega t^{\prime}}e^{2i\varphi}\right)\right]+{\rm c.c.}\right\}\,.$
When $\Omega>0$, the second term oscillates more rapidly than the third term, so the second term can be neglected. When $\Omega<0$, the situation is reversed, and the third term can be neglected. Now we take $\Omega>0$ and performe the integration, we have
$\hskip-12.0ptF^{(+)}(t,r)$
$=$ $\hskip-5.0ptr\int_{0}^{\infty}dk~\frac{1}{2}\int_{0}^{\pi}\sin\theta d\theta~\frac{1}{2\pi}\int_{0}^{2\pi}d\varphi$
$k\left\{\frac{i}{2}e^{ikr\cos\theta-i\omega_{k}t}\left[e_{xx}^{(+)}\sin^{2}\theta\delta(k-0)+e_{xx}^{(+)}\frac{1+\cos^{2}\theta}{2}\delta(k-2\Omega)e^{2i\varphi}\right]+{\rm c.c.}\right\}$
$=$ $\hskip-5.0pt\frac{r~2\Omega}{8\pi}\int_{0}^{\pi}\sin\theta d\theta\int_{0}^{2\pi}d\varphi\left\{ie^{i2\Omega(r\cos\theta-t)}\frac{1}{\sqrt{2}}\left(\cos^{2}\theta\cos^{2}\varphi-\sin^{2}\varphi\right)\frac{1+\cos^{2}\theta}{2}e^{2i\varphi}+{\rm c.c.}\right\}$
$=$ $\hskip-5.0pt\frac{\sqrt{2}}{64(r\Omega)^{4}}\left\{ie^{-i2\Omega t}\left[2r\Omega\left(-3+4(r\Omega)^{2}\right)\cos(2r\Omega)+\left(3-8(r\Omega)^{2}+8(r\Omega)^{4}\sin(2r\Omega)\right)\right]+{\rm c.c.}\right\}\ .$
Since the distance from a binary black holes $r$ is enough larger than the Schwarzschild radius, we can take $r\Omega\gg 1$, and then obtain
$F^{(+)}(t,r)\hskip-5.0pt$ $\simeq$ $\hskip-5.0pt\frac{\sqrt{2}}{64(r\Omega)^{4}}\left\{ie^{i2\Omega t}\left[8(r\Omega)^{4}\sin(2r\Omega)\right]+{\rm c.c.}\right\}$ (A.15)
$=$ $\hskip-5.0pt\frac{1}{4\sqrt{2}}\left[\cos(2\Omega(t+r))-\cos(2\Omega(t-r))\right]\ .$
Similarly, we can calculate $F^{(\times)}(t,r)$ as well.
$F^{(\times)}(t,r)\hskip-5.0pt$ $=$ $\hskip-5.0pt\frac{1}{4\sqrt{2}}\left[\cos(2\Omega(t+r))-\cos(2\Omega(t-r))\right]\ .$ (A.16)
Thus, substituting Eq. (A.15) and (A.16) into Eq. (A.9), we obtain
[…]
$=$ $\hskip-5.0pt\frac{2G\mu(a\Omega)^{2}}{r}\left[\cos(2\Omega(t+r))-\cos(2\Omega(t-r))\right]\,.$
Appendix B Derivation of the squeezing parameter
The squeezing parameter is expressed as
$\hskip-28.45274pt\beta^{(PQ)}_{kk^{\prime}}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{3i}{(2\pi)^{3}2M_{\rm p}^{2}}\sum_{N=1,2}\gamma_{N}^{5}m_{N}\int^{t}dt^{\prime}\frac{e^{i(\omega_{k}+\omega_{k^{\prime}})t^{\prime}}}{2\sqrt{\omega_{k}\omega_{k^{\prime}}}}e^{(P)}_{ij}({\bm{k}})e^{(Q)}_{lm}({\bm{k}}^{\prime})v_{N}^{i}v_{N}^{j}v_{N}^{l}v_{N}^{m}e^{-i\bm{k}\cdot\bar{\bm{x}}_{N}}e^{-i\bm{k}^{\prime}\cdot\bar{\bm{x}}_{N}}\ .$ (B.1)
Since Eq. (B.1) contains the phase factor $\exp\left\{-i\left[(\bm{k}+\bm{k}^{\prime})\cdot\bar{\bm{x}}_{N}-(\omega_{k}+\omega_{k^{\prime}})t^{\prime}\right]\right\}$, by performing the integration with respect to time, we obtain relations $\bm{k}+\bm{k}^{\prime}=0$ and $\omega_{k}=k=2\Omega$. Physically, the first condition reflects conservation of the total momentum of the emitted pair due to translational symmetry about the binary’s center of mass, and the second frequency selection is the same as that given by the classical quadrupole formula. After performing azimuthal integrating of $\bm{k},-\bm{k}$, namely $(\theta^{\prime},\varphi^{\prime})=(\pi-\theta,\varphi+2\pi)$, the cross terms involving ++ and $\times$ modes vanish.
In Eq. (B.1), the kernel involves the tensor contraction
$e^{(P)}_{ij}({\bm{k}})\,e^{(Q)}_{lm}({\bm{k}}^{\prime})\,v_{N}^{i}v_{N}^{j}v_{N}^{l}v_{N}^{m}\,.$
For a given wavevector $\bm{k}$}, the projection of the polarization tensors on the orbital velocity can be written schematically as
[…]
with some functions $F_{+}(\theta)$ and $F_{\times}(\theta)$ of the polar angle $\theta$. Thus the cross term with $P\neq Q$ contains a product of the form
[…]
whose integral over the azimuthal angle $\varphi$ vanishes. More explicitly, we find
$\hskip-12.0pt\int_{0}^{\pi}d\theta\sin\theta\int_{0}^{2\pi}d\varphi\;e^{(+)}_{ij}({\bm{k}})e^{(\times)}_{lm}(-{\bm{k}})v_{1}^{i}v_{1}^{j}v_{1}^{l}v_{1}^{m}\,e^{-i{\bm{k}}\cdot\bar{\bm{x}}_{N}}e^{-i(-{\bm{k}})\cdot\bar{\bm{x}}_{N}}$ (B.2)
$=$ $\frac{1}{2}\left(\frac{m_{2}}{M}a\Omega\right)^{4}\int_{0}^{\pi}d\theta\sin\theta\int_{0}^{2\pi}d\varphi$
$\times\left[\frac{\sin^{2}\theta}{2}+\frac{1+\cos^{2}\theta}{2}\cos(2\Omega t^{\prime}-2\varphi)\right]\cos(\pi-\theta)\,\sin\!\left(2\Omega t^{\prime}-2(\varphi+2\pi)\right)$
$=$ $0\,.$
This shows that, after integration over the azimuthal angles of $\bm{k}$} and $\bm{k}$$^{\prime}$, the cross terms with $P\neq Q$ vanish, and only the diagonal contributions with $P=Q$ survive. By contrast, in the circular (helicity) basis defined by $e_{ij}^{\mathrm{R,L}}=(e_{ij}^{(+)}\pm i\,e_{ij}^{(\times)})/\sqrt{2}$, the kernel generally acquires off–diagonal components because the diagonal components are not equal for our binary configuration. Thus, right- and left-handed polarizations are correlated in this basis.
Therefore, by using Eqs. (A.6) and (A.7), the squeezing parameter can be written explicitly as
$\beta^{(+)}_{kk^{\prime}}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{3i}{64\pi^{3}M_{\rm p}^{2}}\mu(a\Omega)^{4}$ (B.3)
@@ Appendix B Derivation of the squeezing parameter
$\hskip-42.67912pt\times\frac{e^{i(\omega_{k}+\omega_{k^{\prime}})t^{\prime}}}{\sqrt{\omega_{k}\omega_{k^{\prime}}}}\left[\frac{\sin^{2}\theta}{2}+\frac{1+\cos^{2}\theta}{2}\cos(2\Omega t^{\prime}-2\varphi)\right]\left[\frac{\sin^{2}\theta^{\prime}}{2}+\frac{1+\cos^{2}\theta^{\prime}}{2}\cos(2\Omega t^{\prime}-2\varphi^{\prime})\right]\,$,
@@
$\beta^{(\times)}_{kk^{\prime}}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{3i}{64\pi^{3}M_{\rm p}^{2}}\mu(a\Omega)^{4}$ (B.4)
@@
For simplicity, similar to the case of coherent states, we set
$\gamma_{1}^{5}\left(\frac{m_{2}}{M}\right)^{3}e^{-i(k_{x}+k^{\prime}_{x})x_{1}-i(k_{y}+k^{\prime}_{y})y_{1}}+\gamma_{2}^{5}\left(\frac{m_{1}}{M}\right)^{3}e^{-i(k_{x}+k^{\prime}_{x})x_{2}-i(k_{y}+k^{\prime}_{y})y_{2}}\simeq 1\ .$ (B.5)
Based on these discussions, we can write $\beta^{(+)}$ specifically for $\bm{k}=-\bm{k}^{\prime}$, that is, $(\theta^{\prime},\varphi^{\prime})=(\pi-\theta,\varphi+2\pi)$, as follows:
$\beta^{(+)}_{kk^{\prime}}\hskip-5.0pt$ $=$ $\hskip-5.0pt-\frac{3i}{64\pi^{3}M_{\rm p}^{2}}\mu(a\Omega)^{4}\int^{t}dt^{\prime}\frac{e^{i(\omega_{k}+\omega_{k^{\prime}})t^{\prime}}}{\sqrt{\omega_{k}\omega_{k^{\prime}}}}$ (B.6)
$\times\left[\frac{\sin^{2}\theta}{2}+\frac{1+\cos^{2}\theta}{2}\cos(2\Omega t^{\prime}-2\varphi)\right]\left[\frac{\sin^{2}(\pi-\theta)}{2}+\frac{1+\cos^{2}(\pi-\theta)}{2}\cos(2\Omega t^{\prime}-2(\varphi+\pi))\right]$
$=$ $\hskip-5.0pt-\frac{3i}{64\pi^{3}M_{\rm p}^{2}}\mu(a\Omega)^{4}\int^{t}dt^{\prime}\frac{e^{i(\omega_{k}+\omega_{k^{\prime}})t^{\prime}}}{\sqrt{\omega_{k}\omega_{k^{\prime}}}}\left[\frac{\sin^{2}\theta}{2}+\frac{1+\cos^{2}\theta}{2}\cos(2\Omega t^{\prime}-2\varphi)\right]^{2}\,.$
Thus, the magnitude of $\beta$ is estimated
$|\beta|\simeq\frac{3\mu(a\Omega)^{4}}{64\pi^{3}M_{\rm p}^{2}}\frac{1}{k^{2}}\ .$ (B.7)
Note that the strength of the squeeze effect depends on the direction of gravitational wave emission. To determine the squeezing parameter $\zeta$, it is necessary to convert it into a dimensionless and volume independent quantity, so it is weighted by the corresponding wave number space volume $(k=2\Omega)$.
$\zeta\simeq\frac{4\pi}{3}k^{3}|\beta|\simeq\frac{1}{8\pi M_{\rm p}^{2}}\mu(a\Omega)^{4}f\ ,$ (B.8)
where we used the relation $k=2\pi f$.